1.Suppose that airplane engines operate independently in flight and fail with probability 1/5. Assuming that a plane makes a safe flight if at least one-half of its engines run, what is the probability that 4-engine plane has a successful flight?
2.Suppose that airplane engines operate independently in flight and fail with probability 1/5. Assuming that a plane makes a safe flight if at least one-half of its engines run, what is the probability that 2-engine plane has a successful flight?
3. A math teacher observes that on the average 3 students came to their 8-9AM class late. The probability that at most 2 students will be late in their 8-9AM class
4.A math teacher observes that on the average 3 students came to their 8-9AM class late. The probability 5 students will be late in their 8-9AM class is____?

Answer by stanbon(75887)

1.Suppose that airplane engines operate independently in flight and fail with probability 1/5. Assuming that a plane makes a safe flight if at least one-half of its engines run, what is the probability that 4-engine plane has a successful flight?
Binomial Problem with n = 4 and p(fail) = 1/5
Ans: P(0<= x <=2) = binomcdf(4,1/5,2) = 0.9728
--------------------------------------
2.Suppose that airplane engines operate independently in flight and fail with probability 1/5. Assuming that a plane makes a safe flight if at least one-half of its engines run, what is the probability that 2-engine plane has a successful flight?
Ans: P(0<= x <=1) = binomcdf(4,1/4,1) = 0.8192
-------------------------------
3. A math teacher observes that on the average 3 students came to their 8-9AM class late. The probability that at most 2 students will be late in their 8-9AM class
Note: n not defined
Ans: P(0<= x <=2) = binomcdf(n,3/n,2) = ?
4.A math teacher observes that on the average 3 students came to their 8-9AM class late. The probability 5 students will be late in their 8-9AM class is
Note:: n not defined
----------
Cheers,
Stan H.
-------------